# Friederike Becker1, Fabian Krüger1 and Melanie Schienle1,2

### v2

@@ 3.4 Forecast Performance on the Hold-Out Data
Here we use the 2013-2023 hold-out data for evaluating the model configuration ($R=11$, absolute error method) chosen as described above. As noted in Section 3.1, we use the first fall release for defining the truth value against which we evaluate the predictions. [^2] For the year 2023, we use the truth value published in the WEO’s spring 2024 edition, as the fall 2024 release is not yet available at the time of writing this. Furthermore, as the inflation data series that we use for the benchmark models stops for Japan in 2021 and we thus lack benchmark predictions from 2021 onward, Japan is excluded from scoring in the years 2021 to 2023. Figure 3 shows interval scores, whereas Figure 4 shows empirical coverage of the intervals.
@@
For inflation, IMF-based forecasts tend to perform best at the shortest horizon (fall, current year), whereas scores are otherwise similar for all models. The decomposition into the score components is also similar, with dispersion and underprediction accounting for a sizeable portion of the penalties. Empirical coverage values tend to stay close to nominal levels, for all methods.
@@
For GDP growth, the IMF-based forecasts tend to receive better interval scores than competing methods, across all horizons. The score decomposition is similar for next-year forecasts across methods, with overprediction making up the largest portion, followed by dispersion. Recall that the over- and underprediction components scale linearly with the distance of forecast interval and realized value. Thereby, years such as 2020 with pronounced negative growth rates, which generally were not anticipated a year or more in advance, enter particularly heavily into the overprediction component. Concerning calibration, all methods attain close to nominal empirical coverage levels. The IMF-extracted forecasts overall show little variability in empirical coverage across horizons and countries, but are slightly less well calibrated at the $50\%$ level than competing methods.
@@
All in all, we conclude that IMF-extracted forecasts attain similar or sometimes slightly better performance than competing methods, for both targets and main evaluation metrics used.
@@
Figure 5 shows the average length of the prediction intervals of IMF-extracted probabilistic forecasts, separately for each combination of target variable and forecast horizon. For both inflation and real GDP growth, lengths shrink considerably between next-year and same-year horizons, as well as between the two same-year horizons. The reduction in size is less considerable between the two next-year horizons, especially for inflation, suggesting that the average level of confidence stays mostly similar from one-and-a-half years to one year out from the target. This result is in accordance with Qu et al. (2019), who don’t find consistent significant evidence that the WEO’s average forecast accuracy improves during this time period.
@@
(b) Real GDP growth
Figure 3: Interval scores (IS), separately for (a) inflation and (b) real GDP growth, evaluation period 2013-2023. The IS is decomposed into overprediction, underprediction and dispersion, as indicated by the shaded areas. Scores are computed as a weighted average of individual scores at the $50\%$ and $80\%$ level.
@@
Figure 4: Empirical prediction interval coverage levels, for (a) inflation and (b) real GDP growth, evaluation period 2013-2023. In each panel, (I) shows the overall empirical coverage for all methods, with translucent smaller points in the background representing individual coverage for each country-horizon pair, indicating underlying variability in coverage levels. (II) and (III) show coverage levels by forecast horizon, for 50% and 80% nominal coverage, respectively.
@@
Figure 5: Average length of prediction intervals for IMF-extracted probabilistic forecasts, in the years 2013-2023.
3.5 Robustness Checks
Here we describe two robustness checks on potentially important implementation choices of our empirical setup.
@@ 3.5 Robustness Checks
First, we note that the results in Figures 3 and 4 are based on annual truth values provided by the IMF. By contrast, the benchmark models described in Section 3.3 are fitted to higher-frequency (quarterly) data that are not provided by the IMF, and were thus chosen from another source (OECD). While we made our best effort to choose compatible series, full consistency between the two sources seems unrealistic given the complexity of international macroeconomic aggregates. As a consequence, our use of IMF data for evaluation may pose an unfair disadvantage to the benchmark models that were fitted on OECD data. To examine this possibility, we consider an alternative set of annual truth data that we construct from quarterly OECD data. We use these truth data for both the calculation of forecast errors and the subsequent evaluation. Figure 7 in the appendix studies how this change affects interval scores. The figure focuses on the comparison between the IMF-based forecasts and BVAR-direct as an illustrative benchmark method. In a majority of setups (countries, variables and horizons), the different choice of truth value does not change the model ranking between the IMF and BVAR-direct. However, for the shortest horizon (fall, current year), the model ranking is typically reversed in favor of the BVAR-direct method. These results are confirmed by Figures 8 and 9 in the appendix. Thus, while the IMF-based forecasts remain competitive at the other forecast horizons, their good performance at the shortest horizon vanishes when using OECD rather than IMF truth values. While somewhat unsatisfactory, this finding is plausible from a statistical perspective: The shortest forecast horizon is characterized by fairly high predictability since relevant components of the predictand (e.g., GDP growth in 2022) are known when the forecast is issued (e.g., September 2022). In this low-noise setting, the details of the definition of the truth value matter. Hence the IMF-based model outperforms the benchmarks when evaluated against IMF data, whereas the benchmarks prevail when evaluated against OECD data. This effect is not present at longer forecast horizons, where predictability is much lower, so that differences in the definition of the truth value are dominated by other sources of forecast error.
Second, following Adrian et al. (2019), a recent strand of literature discusses the role of financial stress indexes for predicting inflation and GDP growth. To relate to this discussion, we consider augmenting the BVAR benchmark models by the CISS summary indicator of financial stress (Holló et al., 2012). As shown by Figure 10 in the appendix, the BVAR variant including CISS does not systematically outperform the variant without CISS, in terms of interval scores. Thus, our finding that the IMF-based prediction intervals perform competitively remains intact when augmenting the BVAR benchmarks by information on financial stress.
@@ 3.6 Outlook: Prospective Real-Time Forecasts and Evaluation
Initial releases of the forecasts for October 2023 and April 2024 are available in a public GitHub repository (Becker et al., 2023), ensuring transparency and accountability through time stamps. The latter repository also links to the website https://probability-forecasting.shinyapps.io/macropi/ on which we provide concise visualizations of all forecasts and prediction intervals. Moreover, we have deposited a preregistration protocol (Becker et al., 2024) in which we make a commitment regarding how we will evaluate forecasts in 2029, after collecting prospective forecasts and prediction intervals for five years.
@@ 4 Conclusion
This paper presents a method for predicting the distribution of output growth and inflation in the G7 economies. Our method aims to be as simple as possible, subject to being ‘reasonably competitive’, in the sense that we see no straightforward way to improving empirical forecasting performance. While we use point forecasts from the IMF, our suggested procedure can in principle be employed for point forecasts from any source, including macroeconomic model-based forecasts, judgmental forecasts (issued, e.g., by central banks and surveys) or forecasts based on statistical or machine learning models, possibly using a large number of input variables. As an important (and often critical) practical requirement, an informative history of out-of-sample forecast errors must be available in order to estimate the distribution of forecast errors. Our positive empirical results on the IMF forecasts display the common finding that point forecasts by economic institutions are not easy to beat by means of purely statistical models (see e.g. Faust and Wright, 2013).
Appendix
@@ Appendix A Uncertainty Quantification in Central Bank Reports
We are searching for concretely quantified statements about the uncertainty surrounding macroeconomic forecasts in the G7’s central banks’ “flagship” reports for communicating their growth and inflation projections. For this, we applied the following protocol.
@@
Canada Bank of Canada Monetary Policy Report April 2024 no -
France Banque de France Macroeconomic Projections France September 202311footnotemark: 1 no -
Germany Deutsche Bundesbank Monthly Report December 202322footnotemark: 2 no -
Italy Banca d’Italia Macroeconomic Projections for the Italian economy April 2024 no -
Japan Bank of Japan Outlook for Economic Activity and Prices January 2024 somewhat truncated range of board members’ votes33footnotemark: 3
United Kingdom Bank of England Monetary Policy Report February 2024 yes fan chart
United States Federal Reserve Monetary Policy Report April 2024 yes fan chart
Table 3: State of uncertainty quantification for macroeconomic forecasts that are issued by the central banks of the G7 countries. For each central bank, we consulted the “flagship” report containing forecasts for inflation and real GDP growth, often titled “Macroeconomic Projections”, “Economic Outlook” or “Monetary Policy Report” (depending on availability) and checked whether that document’s centrally communicated forecasts are accompanied by an explicit quantification of uncertainty.
1The latest available report as of Mid April 2024.
2Not all monthly reports contain macroeconomic forecasts - the most recent publication containing them was thus chosen.
3These can’t be interpreted as quantiles of the forecast distribution, as every board member is presumably forecasting a central functional (mean, median, etc.) of the forecast distribution.

### v3

@@ 3.4 Forecast Performance on the Hold-Out Data
Here we use the 2013-2023 hold-out data for evaluating the model configuration ($R=11$, absolute error method) chosen as described above. [^3] As the inflation data series that we use for the benchmark models stops for Japan in 2021 and we thus lack benchmark predictions from 2021 onward, Japan is excluded from scoring in the years 2021 to 2023. As noted in Section 3.1, we use the first fall release for defining the truth value against which we evaluate the predictions. Figure 3 shows interval scores, whereas Figure 4 shows empirical coverage of the intervals.
@@
For inflation, IMF-based forecasts tend to perform best at the shortest horizon (“Fall, Current”), whereas scores are otherwise similar for all methods. The decomposition into the score components is also similar, although skewed more towards dispersion for the benchmarks and towards underprediction for IMF-based forecasts, especially at next-year horizons. When considering interval scores separately at the two levels, see Table 4 in the supplement, the IMF-based forecasts are favored at the $50\%$ level, while the benchmarks tend to attain lower score values for the $80\%$ level, especially at the two longer horizons. Empirical coverage values overall tend to stay close to nominal levels, for all methods.
@@
For GDP growth, the IMF-based forecasts tend to receive better interval scores than competing methods, across all horizons. This ordering is also largely upheld when considering interval scores separately at the two levels, see Table 4 in the supplement. The score decomposition is similar for next-year forecasts across methods, with overprediction accounting for the largest portion, followed by dispersion. Recall that the over- and underprediction components scale linearly with the distance of forecast interval and realized value. Thereby, years such as 2020 with pronounced negative growth rates, which generally were not anticipated a year or more in advance, enter particularly heavily into the overprediction component. Concerning calibration, all methods attain close-to-nominal empirical coverage levels. The IMF-based forecasts overall show little variability in empirical coverage across horizons and countries, but are slightly less well calibrated at the $50\%$ level than competing methods.
We also consider Diebold and Mariano (1995) for the null hypothesis of equal predictive accuracy of the IMF-based method and each benchmark method. In order to limit the number of tests conducted, we average the interval scores across countries at each date (for a given variable, forecast horizon and benchmark model). This yields 24 tests in total (two variables, four horizons, and the AR, BVAR and BVAR-direct benchmark models). We consider two-sided tests, a significance level of 95% and use the R package sandwich package and its function NeweyWest (Zeileis, 2004; Zeileis et al., 2020) for the autocorrelation-robust variance estimator that enters the test statistic. Among the 24 test setups, the null of equal predictive accuracy is rejected in favor of the IMF forecasts on four occasions (all referring to GDP growth at long horizons), whereas there are no rejections in favor of a benchmark model. These results partly reflect the short time series dimension of our evaluation sample and the associated low power of the test.
@@
All in all, we conclude that IMF-based forecasts attain similar or sometimes slightly better performance than competing methods, for both targets and main evaluation metrics used.
@@
Figure 5 shows the average length of the prediction intervals of IMF-based probabilistic forecasts, separately for each combination of target variable and forecast horizon. For both inflation and GDP growth, lengths shrink considerably between next-year and same-year horizons, as well as between the two same-year horizons. The reduction in size is less considerable between the two next-year horizons, especially for inflation, suggesting that the average level of confidence stays mostly similar from one-and-a-half years to one year out from the target. This result is in accordance with Qu et al. (2019), who do not find consistent significant evidence that the WEO’s average forecast accuracy improves across these horizons.
@@
(b) GDP growth
Figure 3: Interval scores (IS), separately for (a) inflation and (b) GDP growth, evaluation period 2013-2023. The IS is decomposed into overprediction, underprediction and dispersion, as indicated by the shaded areas. Scores are computed as a weighted sum of interval scores at the $50\%$ and $80\%$ level.
@@
Figure 4: Empirical prediction interval coverage levels, for (a) inflation and (b) GDP growth, evaluation period 2013-2023. In each panel, (I) shows the overall empirical coverage for all methods, with translucent smaller points in the background representing individual coverage for each country-horizon pair, indicating underlying variability in coverage levels. (II) and (III) show coverage levels by forecast horizon, for 50% and 80% nominal coverage, respectively.
@@
Figure 5: Average length of prediction intervals for IMF-based probabilistic forecasts, in the years 2013-2023.
3.5 Further Analyses
Here we summarize further analyses that investigate the robustness of our empirical results (Sections 3.5.13.5.4) and illustrate the properties of our proposed approach (Section 3.5.5).
3.5.1 OECD Truth Values
@@
The results in Figures 3 and 4 are based on annual truth values provided by the IMF. By contrast, the benchmark models described in Section 3.3 are fitted to higher-frequency (quarterly) data that are not provided by the IMF, and were thus chosen from another source (OECD). While we made our best effort to choose compatible series, full consistency between the two sources seems unrealistic given the complexity of international macroeconomic aggregates. As a consequence, our use of IMF data for evaluation may pose an unfair disadvantage to the benchmark models that were fitted on OECD data. To examine this possibility, we consider an alternative set of annual truth data that we construct from quarterly OECD data. We use these truth data for both the calculation of forecast errors and the subsequent evaluation. Figure 9 in the supplement studies how this change affects interval scores. The figure focuses on the comparison between the IMF-based forecasts and BVAR-direct as an illustrative benchmark method. In a majority of setups, the different choice of truth value does not change the model ranking between the IMF and BVAR-direct. However, for the shortest horizon (fall, current year), the model ranking is typically reversed in favor of the BVAR-direct method. These results are confirmed by Figures 10 and 11 in the supplement. Thus, while the IMF-based forecasts remain competitive at the other forecast horizons, their good performance at the shortest horizon vanishes when using OECD rather than IMF truth values. While somewhat unsatisfactory, this finding is plausible from a statistical perspective: The shortest forecast horizon is characterized by fairly high predictability since relevant components of the predictand (e.g., GDP growth in 2022) are known when the forecast is issued (e.g., September 2022). In this low-noise setting, the details of the definition of the truth value matter. Hence the IMF-based model outperforms the benchmarks when evaluated against IMF data, whereas the benchmarks prevail when evaluated against OECD data. This effect is not present at longer forecast horizons, where predictability is much lower, so that differences in the definition of the truth value are dominated by other sources of forecast error.
3.5.2 Additional Intervals and Measures of Forecast Performance
In the main part of the paper, our analysis of forecast performance has focused on prediction intervals at the 50% and 80% level, i.e., $\tau\in\{0.5,0.8\}$, which we consider particularly relevant in practical applications. Here, we consider an extension where we generate prediction intervals at equally spaced levels $\tau\in\{0.1,0.2,...,0.9\}$, and additionally report more evaluation measures, to investigate robustness of results with respect to these extensions.
In Tables 6 and 7 in the supplement, we report the evaluation measures for these forecast intervals, including various types of interval scores as well as the Continuous Ranked Probability Score (CRPS; Matheson and Winkler 1976) that each consider the entire predictive distribution. In addition, some individual interval scores at representative levels $\tau\in\{0.3,0.5,0.8,0.9\}$ are shown in Table 8 in the supplement. These measures yield the same qualitative conclusions as our previous analysis.
Additionally, Figure 12 in the supplement presents coverage statistics for the corresponding forecast quantiles, where a forecast interval with $\alpha=1-\tau$ induces the forecast quantiles $\hat{q}_{\alpha/2}$ and $\hat{q}_{1-\alpha/2}$. As shown there, the proposed IMF-based method attains close-to-nominal calibration in most instances, whereas some of the benchmark models exhibit considerable deviations from nominal coverage, especially for GDP growth. Nevertheless, for inflation and quantile levels larger than 0.85, some benchmarks, e.g., BVAR-direct, attain better calibration than the IMF-based forecasts.
3.5.3 Additional Benchmark Models
In addition to the AR(1) and BVAR models described earlier, Tables 6 and 7 in the supplement present results for additional benchmark models. Results are shown for both the extended set of intervals introduced in the previous section and the base set of intervals from the main analysis. Conceptually, these models are variations and extensions of the main benchmarks, to investigate robustness of results with respect to concrete benchmark specifications. These include an AR($p$) model, where the lag length parameter $p$ is selected recursively using the Schwarz (1978) information criterion; the ‘BVAR-CISS’ model including a summary indicator of financial stress (Holló et al., 2012); the ‘BVAR-Mix’ model that uses stochastic volatility prior to 2019 and constant volatility afterwards (see Section 3.3); and an ensemble using equally weighted quantile-wise averages of the IMF, AR and BVAR forecasts. The ensemble is a natural benchmark given extensive empirical evidence on the good performance of forecast combinations in various empirical setups (Wang et al., 2023). In addition to benchmarks trained on quarterly OECD data, we fit two autoregressive specifications on annual outcome data provided by the IMF, and thereby tailor models directly to the annual growth rates to be predicted. While ‘AR-annual’ only has access to the latest annual truth value, ‘ARX-annual’ compensates for this informational disadvantage by incorporating the most recent IMF point forecasts as regressors. The latter is thus based on the same information set as our proposed approach. Section B.5 in the supplement provides more detailed descriptions of the additional benchmarks.
We find that previous results are in general robust to concrete specifications of the benchmark models. ‘AR-annual’ exhibits high scores on current-year forecasts, where its informational disadvantage is particularly high, but performs comparably to other models on next-year forecasts. In terms of the CRPS, some alternative specifications outperform the respective counterparts from the main analysis on a subset of target-horizon combinations. In particular, the AR($p$) specification outperforms the AR(1) variant for inflation at the next-year horizons. Overall, however, any improvements of alternative specifications over the main benchmarks tend to be small and inconsistent across setups. Compared to our proposed IMF-based approach, the BVAR-Mix model performs slightly better for inflation at both next-year horizons, but performs similar or worse otherwise. This model additionally requires a manual choice of regime switch (necessitated by Covid-19 outliers) which can be viewed as a conceptual drawback as compared to the other models we consider. The ARX-annual method performs worse than the IMF-based approach in most setups. Given that both methods are based on the same information set, this implies that the modeling assumptions of the IMF-based approach are more suitable from a forecasting perspective. Among all benchmarks, the ensemble tends to perform best overall, and similarly well as our proposed IMF-based method.
3.5.4 Additional Countries
An attractive feature of our methodology is its flexibility to be readily extended to generate prediction intervals for additional countries. This is particularly appealing in conjunction with the IMF WEO point forecasts, which have nearly worldwide coverage.
Thus, while our focus remains on the G7 economies, we extend a version of our analysis to ten additional countries in order to investigate the robustness of our approach. The set of benchmark models here is reduced to those that are fit directly on annual IMF truth data. We choose each of the five largest Emerging Market Economies and Advanced Economies (besides the G7), with data on country classification and economy size obtained from the IMF (International Monetary Fund, 2024b). These criteria yield the following countries: China, India, Brazil, Russia and Mexico (Emerging Market Economies) and South Korea, Australia, Spain, the Netherlands and Switzerland (Advanced Economies). We thereby keep focus on a comparable number of countries to the originally considered G7, while still retaining some generality across different economy types.
As shown in Figures 13 and 14 in the supplement, our proposed approach attains close-to-nominal coverage for the extended set of countries as well. While aggregate coverage levels are in some cases slightly better for the AR-annual and ARX-annual models, the IMF-based forecasts exhibit lower levels of variability in empirical interval coverage across countries. Variability across horizons is similarly small for all methods. In terms of different variants of the interval score, see Tables 9 and 10 in the supplement, our proposed method consistently outperforms both benchmarks, and in particular the ARX-annual method that is based on the same information set.
3.5.5 Forecasts During the Covid-19 Period
Tables 11 and 12 in the supplement present our main performance measures for the period 2020–2023, that is, the Covid-19 pandemic period and its immediate aftermath. We present results for this sub-period in order to investigate whether or not our results are driven by a small number of extreme observations. That said, this analysis should be viewed as illustrative since conditioning on extreme events is subject to the “forecaster’s dilemma” critique (Lerch et al., 2017) and is thus not helpful in terms of formal forecast evaluation.
For GDP growth, the IMF-based forecasts continue to outperform the benchmarks in terms of different variants of the interval score, for three out of four horizons. For inflation, the IMF-based forecasts perform worse than the benchmarks in all but the shortest horizon. In conjunction with the interval score results for the entire hold-out sample (2013–2023, see Figure 3, and additionally Tables 6 and 7 in the supplement), this finding indicates that our proposed IMF-based method generally outperforms the benchmarks for the 2013–2019 period, whereas the ranking for the Covid-19 period is mixed as described. Unsurprisingly, all methods attain less–than-nominal quantile coverage rates during the Covid-19 period (on average across years and countries). That said, the coverage rates of the IMF-based method are clearly lower than for the benchmark methods.
Furthermore, the Covid-19 period is interesting to showcase how our proposed method’s prediction intervals are updated over time. Compared to some of the benchmark models, our method’s updating process is rather insensitive to shocks, which is due to two main features. First, the use of quantiles, which clearly limits the impact of extreme observations. Second, the use of annual (as opposed to quarterly) data on forecast errors. Figures 15, 16, and 17 in the supplement illustrate this point by showing prediction intervals before and after years with large forecast errors (2020 for GDP growth and 2022 for inflation).
From an ex post perspective, low responsiveness to Covid-related shocks turned out advantageous in the case of GDP growth, where the extreme forecast error in 2020 was followed by moderate forecast errors afterwards. For inflation, where IMF-based intervals were already quite narrow compared to the benchmark models, the low immediate reactiveness of the quantile-based method was a disadvantage, especially if the increase in forecast errors after 2022 represents a continued structural change beyond the end of our study period, which ends in 2023. Within the scope of our prospective study, see Section 3.6, we will be able to monitor the further performance of IMF-based forecast intervals for inflation.
Apart from statistical forecasting performance, our method’s use of quantiles and annual data clearly contributes to its simplicity and transparency. By contrast, the forecasts and updating mechanisms of statistical and machine learning methods tend to be far more complicated and may depend on hyperparameters in non-obvious ways.
@@ 3.6 Outlook: Prospective Real-Time Forecasts and Evaluation
Initial releases of the forecasts for October 2023 and April 2024, and October 2024 are available in a public GitHub repository (Becker et al., 2023), ensuring transparency and accountability through time stamps. The latter repository also links to the website https://probability-forecasting.shinyapps.io/macropi/ on which we provide concise visualizations of all forecasts and prediction intervals. Moreover, we have deposited a preregistration protocol (Becker et al., 2024) that covers forecasts from October 2024 onwards. In this protocol, we make a commitment regarding how we will evaluate forecasts in 2029, after collecting prospective forecast intervals for five years.
@@ 4 Conclusion
This paper presents a method for predicting the distribution of output growth and inflation in the G7 economies. Our method aims to be as simple as possible, subject to being “reasonably competitive”, in the sense that we see no straightforward way to improving empirical forecasting performance. In particular, assessing the impact of our method’s tuning parameters (the number of past forecast errors considered, as well as the use of absolute versus raw forecast errors) is easier than in time series models that achieve comparable forecast performance. Among the benchmark methods we consider, an ensemble of our proposed approach with two time series models performs similarly well in terms of overall forecast accuracy. At the same time, the ensemble is harder to replicate and communicate in that obtaining and explaining its forecasts requires input from all three component forecasts. From the perspective of forecast performance, potential limitations of our quantile-based method include its limited consideration of extreme events and a comparably slow adaptation to potential structural change in the variability of forecast errors. This behavior, however, has also led to some robustness of results e.g. during the Covid-19 period but will be more closely examined in our prospective study.
While we use point forecasts from the IMF, our suggested procedure can in principle be employed for point forecasts from any source, including macroeconomic model-based forecasts, judgmental forecasts (issued, e.g., by central banks and surveys), forecasts based on statistical or machine learning models, possibly using a large number of input variables. This versatility is useful in practice, where point forecasts are often readily available, be it via existing data sources (as for macroeconomic survey forecasts) or software packages (as for model based forecasts). As an important and often critical practical requirement, an informative history of out-of-sample forecast errors must be available in order to estimate the distribution of forecast errors. From an economic forecasting perspective, our empirical results on the IMF forecasts display the common finding that point forecasts by economic institutions are not easy to beat by means of purely statistical models (see e.g. Faust and Wright, 2013).
Supplement
@@ Appendix A Uncertainty Quantification in Central Bank Reports
We are searching for concretely quantified statements about the uncertainty surrounding macro-economic forecasts in the G7’s central banks’ “flagship” reports for communicating their growth and inflation projections. For this, we applied the following protocol.
@@
Canada Bank of Canada Monetary Policy Report January 2025 no -
France Banque de France Macroeconomic Projections France December 2024 no -
Germany Deutsche Bundesbank Monthly Report December 202411footnotemark: 1 no -
Italy Banca d’Italia Macroeconomic Projections for the Italian Economy April 2025 yes fan chart
Japan Bank of Japan Outlook for Economic Activity and Prices January 2025 somewhat truncated range of board members’ votes22footnotemark: 2
United Kingdom Bank of England Monetary Policy Report February 2025 yes fan chart
United States Federal Reserve Monetary Policy Report February 2025 yes multiple, including fan chart
Table 3: State of uncertainty quantification for macroeconomic forecasts that are issued by the central banks of the G7 countries. For each central bank, we consulted the latest available “flagship” report containing forecasts for inflation and real GDP growth, often titled “Macroeconomic Projections”, “Economic Outlook” or “Monetary Policy Report” (depending on availability) and checked whether that document’s centrally communicated forecasts are accompanied by an explicit quantification of uncertainty. All reports were retrieved from the respective bank’s website on April 14, 2025.
1Not all monthly reports contain macroeconomic forecasts - the most recent publication containing them was thus chosen.
2These can’t be interpreted as quantiles of the forecast distribution, as every board member is presumably forecasting a central functional (mean, median, etc.) of the forecast distribution.